Errata for “Noncompact simplexes in Banach spaces with the Radon-Nikodym property” Richard D. Bourgin & G. A. Edgar

Page 164 Line -3. Replace ∑_{j=1} b_{j} by ∑_{j=1}^{k} b_{j} Page 166 Line -8. Delete the entire parenthesis Page 167 Line 3. After 𝒢 Add , f_{1} is 𝒢-measurable

Page 169 Line 15. Replace ∫_{G×A} h dγ by ∫_{G×A} h dν

Page 171 Line -8. Replace g_{n+1}(i_{1}, …, i_{n}, 0) by g_{n+1}(i_{1}, …, i_{n}, 0, —) Page 171 Line -7. Replace g_{n+1}(i_{1}, …, i_{n}, 1) by g_{n+1}(i_{1}, …, i_{n}, 1, —) Page 171 Line -5. Replace g_{n+1}(i_{1}, …, i_{n+1}) by g_{n+1}(i_{1}, …, i_{n+1}, —) Page 171 Line -3. Replace f_{n+1}(j_{1}, …, j_{n+1}) by f_{n+1}(j_{1}, …, j_{n+1}, —) Page 173 Line -14. Replace [4, Theorem 5.1] by [6, Theorem 5.1]

Typeset by AMS-TEX